OpenStudy (anonymous):

Systems of Equations, please help =) Solve without matrix 5x-y+z=4; x+2y-2=5; 2x+3y-3=5 I have the answer, but have tried multiple time and cannot get it.

zepdrix (zepdrix):

System of three equations? :( Oh boy these are not fun lol

zepdrix (zepdrix):

Oh two of them only involve x and y! yay! :)

zepdrix (zepdrix):

Start there, with the x/y equations, find a solution for x or y.

zepdrix (zepdrix):

\[\large\rm x+2y-2=5\]\[\large\rm 2x+3y-3=5\]Let's get these into `standard form`: In which only x's and y's appear on the left sides of the equations.

zepdrix (zepdrix):

So we'll add 2, and 3, respectively, ya?\[\large\rm x+2y=7\]\[\large\rm 2x+3y=8\]Ok with that step? :o

OpenStudy (anonymous):

Yes. Thank you! So sorry I got distracted.